OpenStudy (jaweria):

Can anyone please help me with my Calculus questions? I will appreciate your help!!

OpenStudy (campbell_st):

whats the question...?

OpenStudy (jaweria):

oh thanks there is someone who can help me :)

OpenStudy (campbell_st):

perhaps... depends on the question

OpenStudy (jaweria):

we are looking for f(x) when f '(x) is giving

OpenStudy (jaweria):

f ' (x) = (x+3)(2x-5)

OpenStudy (campbell_st):

ok... so its integral calculus... distribute f'(x) so you have f'(x) = 2x^2 + x - 15 now it should be easier to find the indefinite integral.

OpenStudy (jaweria):

if you dont mind can you help me out doing this question bcoz my teacher didnt explain it well.

OpenStudy (campbell_st):

ok... the the basics of integration is \[\int\limits x^n dx = \frac{1}{n + 1}x^{n + 1} + c\] you need to apply this rule to each term so start with \[2x^2\] add 1 to the power.. and write 1 over the new power.

OpenStudy (jaweria):

ok

OpenStudy (campbell_st):

here is the 1st term \[\int\limits2x dx = 2 \times \frac{1}{3}x^3 = \frac{2}{3}x^3\]

OpenStudy (jaweria):

oh ok

OpenStudy (campbell_st):

oops should have read \[\int\limits 2x^2 dx = 2 \times \frac{1}{3} x^3 ..= \frac{2}{3}x^3\]

OpenStudy (jaweria):

ok

OpenStudy (jaweria):

so thats 2 times where is 2x

OpenStudy (campbell_st):

well its 2 times 1/(x+1) x^(n +1)

OpenStudy (jaweria):

ok

OpenStudy (jaweria):

so what will be our next step?

OpenStudy (campbell_st):

well you need \[\int\limits x.. dx \] what do you think that will be..?

OpenStudy (jaweria):

hmmm I m really not getting this :(

OpenStudy (campbell_st):

do what would you differentiate to get x..?

OpenStudy (jaweria):

x^3?

OpenStudy (jaweria):

I m stuck can you correct me ?

OpenStudy (campbell_st):

not quite if you differentiated and get x, it means you started with \[\frac{x^2}{2}\]